OpenStudy (lolololhelpme):

When traveling a set distance, the speed you travel varies inversely with the time you are traveling. The first time you drive to the store from your house you travel an average of 35 mph and it takes you 10 minutes to get there. The next time you go during school zone hours, so your average speed decreases to 20 mph. How long will it take you to get to the store this time?

OpenStudy (lolololhelpme):

A. 15.5 min B. 17.5 min C. 21 min D. 9 min

OpenStudy (wckeller):

In order to solve this problem, we need to find the distance we are traveling. Do you know the distance formula?

OpenStudy (lolololhelpme):

No I don't think so

OpenStudy (wckeller):

\[speed = \frac{ distance }{ time }\] \[distance = speed*time\]

OpenStudy (wckeller):

First, we need to convert minutes into hours, as our speed is in miles per hour. \[\frac{ 10\min }{ 60\min } =\frac{1}{6}th\] of an hour.

OpenStudy (wckeller):

I think I have made a mistake... I've really confused myself...

OpenStudy (wckeller):

give me a little bit

OpenStudy (lolololhelpme):

lol yeah i'm confused alittle too

OpenStudy (wckeller):

speed = 35 mph time = 1/6 hrs distance=speed*time 35*(1/6)=5.833miles time=distance/speed 5.833miles/20mph = 0.29165 hrs 0.29165hrs * 60min = 17.499 minutes Rounded up = 17.5 min Final answer = B

OpenStudy (wckeller):

I was getting confused because the answer was coming out in hours, I didn't realize I needed to convert it to minutes.

OpenStudy (wckeller):

Sorry it took so long!! I really confused myself

OpenStudy (lolololhelpme):

it's okay!! Thank you!!!

OpenStudy (wckeller):

No problem :)